OpenStudy (anonymous):

how do i graph .21x + .23y Less than or equal to 75

OpenStudy (raden):

first, let that's an equation of straight line :  with (0,a) is y-intercept and (0,b) is x-intercept

OpenStudy (raden):

can u determine the values of a and b resfectively ?

OpenStudy (anonymous):

isnt it .21 and .23?

OpenStudy (raden):

no, to get the y-intercept take x = 0 then solve for y u'll get the value of a while to get the x-intercept take y = 0 then solve for x u'll get the value of b

OpenStudy (raden):

.21x + .23y = 75 if x = 0 then .021(0) +.23y = 75 .23y = 75 y = 75/.23 = 7500/23 so, as y-intercept at point (0, 7500/23)

OpenStudy (e.mccormick):

What RadEn is saying is perfectly correct. Here is just another way to think about it. First, solve for y. That gets it to the line equation you may be more familliar with and you can do what RadEn is talking about. You graph the line because this is \(\le\). If it was just < or >, the line would be dashed because it is not part of the answer. Then shade the proper side of it. However, be careful about the rules for solving inequalities. If you multiply or divide by a negative, the direction of the inequality changes!

OpenStudy (anonymous):

it should be shaded above?

OpenStudy (e.mccormick):

Depends on what you get for an equation, if x>y or x<y. If x>y, below. And x<y, above.

OpenStudy (e.mccormick):

If you leave it as x+y, it can be hard to see the relationship. That is why I like to solve for y first.

OpenStudy (anonymous):

okay but it started at less than or equal to now im confused at what way i shade

OpenStudy (e.mccormick):

\(.21x + .23y\le75\) \(.21x + .23y-.32y\le75-.32y\) \(.21x \le75-.32y\) \(.21x-75 \le75-75-.32y\) \(.21x-75 \le-.32y\) \((.21x-75)/(-.32) \ge-.32y/(-.32)\) \((.21x-75)/(-.32) \ge y\)

OpenStudy (e.mccormick):

Do you see where I divided by a negative and the direction changed?

OpenStudy (anonymous):

yes

OpenStudy (e.mccormick):

That final formula is what I would graph.

OpenStudy (raden):

alternative : take a point (almost it is (0,0)), then subtitute to system inequality above : .21x + .23y <= 75 .21(0,0) + .23(0,0) <= 75 0 <= 75 this is true, so the shaded area would forward to point (0,0). in other words it is under that line

OpenStudy (anonymous):

Thank you!

OpenStudy (e.mccormick):

RadEn has another great point there! Testing parts of the graph aove and below the line for proper results is a great idea. Can clear up where to shade in seconds.

OpenStudy (e.mccormick):

Once you finish your graph, here is one for comparison. Whichever way you like, RadEn's or mine, is fine. They both work and that is the thing about math. He is going right for intercepts and the line, then finishing with a test point. I am putting the line in a format I am more comfortable with and using my understanding to decide where to shade after that. Everybody needs to understand math the way that works for them and get valid answer so they can use it properly.

OpenStudy (anonymous):

thank you!